# Dissipative solutions to Stochastic 3D Euler equations

### v2

@@ 5.1. Estimates on the current
where the first inequality is due to $\mathcal{R}$, the second to a trivial adaptation of [DLK23, Lemma A.4], ˜A.5 and $(l\lambda_{q+1})^{-h_{0}}\lambda^{2}_{q+1}\leq l$. By a similar reasoning as above, we can estimate $\|D_{t,l}(\mathcal{R}\mathbf{P}_{\geq\lambda_{q+1}}G_{i})\|_{M}$. Indeed, it holds:
@@
Thanks to the previous inequality, Eqs.˜3.64 and 4.32, we have that
@@
Therefore, Eqs.˜5.95, 5.100 and 5.102 imply that
(5.103) $\|D_{t,l}\mathcal{R}\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}\lesssim\lambda^{N-1}_{q+1}\frac{\rho}{\mu}\left(\frac{\lambda_{q}\delta_{q}^{1/2}\epsilon^{-1}}{\lambda_{q+1}^{2}}+\frac{\lambda_{q}^{2}\delta_{q}\,l}{\lambda_{q+1}}\right).$
Instead, Bernstein inequality, Eqs.˜5.93 and A.60 imply that
(5.104) $\|D_{t,l}\mathcal{R}(\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}:\nabla u_{l}^{T})\|_{N-1}$ $\lesssim\lambda_{q+1}^{N-1}\big[\|D_{t,l}(\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}:\nabla u_{l}^{T})\|_{0}$
(5.105) $+\|[u_{l}\cdot\nabla,\mathcal{R}](\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}:\nabla u_{l}^{T})\|_{0}\big]$
(5.106) $\lesssim\lambda_{q+1}^{N-1}\Big[\big(\|D_{t,l}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}\|_{0}+\|[u_{l}\cdot\nabla,\mathcal{R}](\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i})\|_{0}\big)\|u_{l}\|_{1}$
@@
At last, Eqs.˜3.51 and 5.84
@@
This concludes the estimates for $\phi_{D_{2}}$. Regarding the estimates on $f_{D_{2}}$, notice that
(5.110) $\int_{\mathbb{T}^{3}}\left(\mathcal{R}(G_{i}):\nabla u_{l}^{T}\right)(x,s)\,dx=\int_{\mathbb{T}^{3}}\left(\mathcal{R}(\mathbf{P}_{<3\lambda_{q+1}/8}G_{i}):\nabla u_{l}^{T}\right)(x,s)\,dx.$
@@
5.1.3. Proof of Proposition˜5.1: bounds on $\phi_{S}$
@@
where the second inequality we have used Lemmas˜3.4, 3.60, 4.11 and 4.40 and
@@
Finally, upon recalling the definition (4.3) of the pressure increment $\overline{p}$, the remaining term $\tilde{\phi}_{S_{1}}:=(\overline{z}\otimes\overline{z}+u_{q}\odot\overline{z})v_{l}+\overline{p}w$ is controlled by means of Proposition˜3.3, Lemma˜3.4, the bounds (2.14), (3.60), (3.51), (4.38), (3.61), and the following inequality:
@@
where the last inequality is due to Lemma˜3.4 and the estimates (4.11) and (4.12).
The next term we study is $\tilde{\phi}_{S_{2}}:=\overline{\mathcal{R}}[(u_{q}\,\mathring{\odot}\,\overline{z}+\overline{z}\otimes\overline{z}):\nabla z_{q}^{T}]$, which is controlled similarly to bounds (5.122) and (5.123) previously obtained for $\tilde{\phi}_{S_{1}}$, with the difference that (5.121) is now replaced by the bound $\|D_{t,q+1}\nabla z_{q}\|_{N}\lesssim i_{q}^{-1/2-\delta}+\begin{cases}1,\quad&N=0\\ \lambda_{q+1}^{N}\delta_{q+1}^{1/2},\quad&N>0\end{cases}$, consequence of Lemma˜3.4, giving
@@
At last, Lemma˜3.4, (4.38), and the following inequalities
@@
5.1.4. Proof of Proposition˜5.1: Conclusion
@@
Bounds (5.34) and (5.35) follow from Lemma˜5.2, Lemma˜5.3, Lemma˜5.4, Lemma˜5.5, Lemma˜5.6 above and our choice of parameters, up to taking the constant $a_{0}>1$ large enough so to absorb all the implicit constants coming from the aforementioned lemmas. As for the estimates on $f$, we have $f:=f_{A}+f_{D_{2}}+f_{S_{2}}$ and thus by Lemma˜5.2, Lemma˜5.4, and Lemma˜5.6:
(5.139) $|\partial_{t}f(t)|\leq|\partial_{t}f_{A}(t)+\partial_{t}f_{D_{2}}(t)+\partial_{t}f_{S_{2}}(t)|\lesssim\lambda_{q+1}^{-3\gamma/2}\delta_{q+2}\delta_{q+1}^{1/2}.$

### v3

@@ 5.1. Estimates on the current
where the first inequality is due to $\mathcal{R}$, the second to a trivial adaptation of [DLK23, Lemma A.4], A.1 and $(l\lambda_{q+1})^{-h_{0}}\lambda^{2}_{q+1}\leq l$. By a similar reasoning as above, we can estimate $\|D_{t,l}(\mathcal{R}\mathbf{P}_{\geq\lambda_{q+1}}G_{i})\|_{M}$. Indeed, it holds:
@@
Thanks to the previous inequality, (3.63), (4.32), we have that
@@
Therefore, (5.95), (5.100), (5.102) imply that
(5.103) $\|D_{t,l}\overline{\mathcal{R}}\mathbf{P}_{\geq\lambda_{q+1}}H_{i}\|_{N-1}\lesssim\lambda^{N-1}_{q+1}\frac{\rho}{\mu}\left(\frac{\lambda_{q}\delta_{q}^{1/2}\epsilon^{-1}}{\lambda_{q+1}^{2}}+\frac{\lambda_{q}^{2}\delta_{q}\,l}{\lambda_{q+1}}\right).$
Instead, Bernstein inequality, (5.93), (A.61) imply that
(5.104) $\|D_{t,l}\overline{\mathcal{R}}(\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}:\nabla u_{l}^{T})\|_{N-1}$ $\lesssim\lambda_{q+1}^{N-1}\big[\|D_{t,l}(\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}:\nabla u_{l}^{T})\|_{0}$
(5.105) $+\|[u_{l}\cdot\nabla,\overline{\mathcal{R}}](\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}:\nabla u_{l}^{T})\|_{0}\big]$
(5.106) $\lesssim\lambda_{q+1}^{N-1}\Big[\big(\|D_{t,l}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}\|_{0}+\|[u_{l}\cdot\nabla,\overline{\mathcal{R}}](\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i})\|_{0}\big)\|u_{l}\|_{1}$
@@
For later use, we additionally notice that the above implies in particular
(5.109) $\|D_{t,l}(\mathcal{R}\mathbf{P}_{\lesssim\lambda_{q+1}}G_{i}:\nabla u_{l}^{T})\|_{0}\lesssim\frac{\rho}{\mu}(\lambda_{q+1}\mu)^{-n_{0}-1}\left[\,\lambda_{q+1}\lambda_{q}\delta_{q}^{1/2}\right],\quad\forall i\in\{1,2,3\}.$
At last, (3.50), (5.84)
@@
This concludes the estimates for $\phi_{D_{2}}$. Regarding the estimates on $f_{D_{2}}$, notice that for every $s\leq\mathfrak{t}$
(5.111) $\int_{\mathbb{T}^{3}}\left(\mathcal{R}(G_{i}):\nabla u_{l}^{T}\right)(s,x)\,dx=\int_{\mathbb{T}^{3}}\left(\mathcal{R}(\mathbf{P}_{<3\lambda_{q+1}/8}G_{i}):\nabla u_{l}^{T}\right)(s,x)\,dx.$
@@
(5.112) $|f_{D_{2}}|+|f^{\prime}_{D_{2}}|\lesssim\frac{\rho}{\mu}\lambda_{q}\delta_{q}^{1/2}(\lambda_{q+1}\mu)^{-n_{0}-1}.$
5.1.3. Proof of Proposition 5.1: bounds on $\phi_{S}$
@@
where the second inequality we have used Lemma 3.3, (3.59), (4.11), (4.40) and
@@
Finally, upon recalling the definition (4.3) of the pressure increment $\overline{p}$, the remaining term $\tilde{\phi}_{S_{1}}:=(\overline{z}\otimes\overline{z}+u_{q}\odot\overline{z})v_{l}+\overline{p}w$ is controlled by means of Proposition 3.1, Lemma 3.3, the bounds (2.14), (3.59), (3.50), (4.38), (3.60), and the following inequality:
@@
where the last inequality is due to Lemma 3.3 and the estimates (4.11) and (4.12).
The next term we study is $\tilde{\phi}_{S_{2}}:=\overline{\mathcal{R}}[(u_{q}\,\mathring{\odot}\,\overline{z}+\overline{z}\otimes\overline{z}):\nabla z_{q}^{T}]$, which is controlled similarly to bounds (5.123) and (5.124) previously obtained for $\tilde{\phi}_{S_{1}}$, with the difference that (5.122) is now replaced by the bound $\|D_{t,q+1}\nabla z_{q}\|_{N}\lesssim i_{q}^{-1/2-\delta}+\begin{cases}1,\quad&N=0\\ \lambda_{q+1}^{N}\delta_{q+1}^{1/2},\quad&N>0\end{cases}$, consequence of Lemma 3.3, giving
@@
At last, Lemma 3.3, (4.38), and the following inequalities
@@
5.1.4. Proof of Proposition 5.1: Conclusion
@@
Bounds (5.34) and (5.35) follow from Lemma 5.1, Lemma 5.2, Lemma 5.3, Lemma 5.4, Lemma 5.5 above and our choice of parameters, up to taking the constant $a_{0}>1$ large enough so to absorb all the implicit constants coming from the aforementioned lemmas. As for the estimates on $f$, we have $f:=f_{A}+f_{D_{2}}+f_{S_{2}}$ and thus by Lemma 5.1, Lemma 5.3, and Lemma 5.5:
(5.140) $|\partial_{t}f|\leq|\partial_{t}f_{A}+\partial_{t}f_{D_{2}}+\partial_{t}f_{S_{2}}|\lesssim\lambda_{q+1}^{-3\gamma/2}\delta_{q+2}\delta_{q+1}^{1/2}.$